# Implicit Bias of Mirror Descent for Shallow Neural Networks in Univariate Regression

### v1

@@ A.1 Linearization
In the following, we use the notation $\Delta=\sup_{t\geq 0}\|H(t)-\tilde{H}(t)\|_{2}$. Now we show $\mathbf{f}(t)$ remains close to $\tilde{\mathbf{f}}(t)$. Let $\mathbf{r}(t)=\mathbf{f}(t)-\tilde{\mathbf{f}}(t)$ and $C_{2}$ be the constant in Proposition 13. Consider the function $u(t)=\|\mathbf{r}(t)\|_{2}$ defined on $t\geq 0$. Clearly, $u(t)$ is continuous. Since $u(t)=0$ if and only if $\mathbf{r}(t)=0$, we have that $u(t)$ is differentiable whenever $u(t)\neq 0$. Note by Proposition 13, we have that, with high probability, $\lambda_{\min}(H(t))\geq C_{2}$ for all $t\geq 0$. Then when $u(t)\neq 0$, we have
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))^{2}$ $=\frac{\mathrm{d}}{\mathrm{dt}}\|e^{C_{2}\eta_{0}t}\mathbf{r}(t)\|_{2}^{2}$ (33)
$=2e^{C_{2}\eta_{0}t}\mathbf{r}(t)^{T}\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}\mathbf{r}(t))$
$=2e^{2C_{2}\eta_{0}t}\mathbf{r}(t)^{T}\Big{(}C_{2}\eta_{0}\mathbf{r}(t)-\eta_{0}H(t)(\mathbf{f}(t)-\mathbf{y})+\eta_{0}\tilde{H}(t)(\tilde{\mathbf{f}}(t)-\mathbf{y})\Big{)}$
$=2\eta_{0}e^{2C_{2}\eta_{0}t}\Big{(}\mathbf{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\mathbf{f}}(t)-\mathbf{y})-\mathbf{r}(t)^{T}(H(t)-C_{2}I_{n})\mathbf{r}(t)\Big{)}$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}\mathbf{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\mathbf{f}}(t)-\mathbf{y})$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}u(t)\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2}.$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))\leq\eta_{0}e^{C_{2}\eta_{0}t}\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\mathbf{f}}(t)-\mathbf{y}\|_{2}.$
Now for any $t\geq 0$, consider $t^{\prime}=\inf\{s\in[0,t]\colon\mathbf{r}(s)\neq 0\}$. Then we have
$e^{C_{2}\eta_{0}t}u(t)\leq\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s.$
@@
$\|\mathbf{r}(t)\|_{2}$ $\leq e^{-C_{2}\eta_{0}t}\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s$ (34)
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}\Delta\int_{t^{\prime}}^{t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}\mathrm{d}s$
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}\|\mathbf{f}(0)-\mathbf{y}\|_{2}t\Delta.$
Given $x\in\mathbb{R}$, by the chain rule we have:
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\theta(t))$ $=-\eta_{0}H(t,x)(\mathbf{f}(t)-\mathbf{y}),$ (35)
where $\text{where }H(t,x)$ $\triangleq\frac{1}{n}J_{\theta}f(x,\theta(t))(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}\mathbf{f}(t)^{T}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\tilde{\theta}(t))$ $=-\eta_{0}\tilde{H}(t,x)(\tilde{\mathbf{f}}(t)-\mathbf{y}),$ (36)
where $\text{where }\tilde{H}(t,x)$ $\triangleq\frac{1}{n}J_{\boldsymbol{a}}f(x,\hat{\theta})(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\boldsymbol{a}}\mathbf{f}(0)^{T}.$
@@
$\|J_{\theta}\mathbf{f}(0)\|_{F}^{2}$ $=\sum_{i=1}^{m}\Big{(}1+\sum_{k=1}^{n}|\sigma(w_{k}(0)x_{i}-b_{k}(0))|^{2}$ (38)
$\quad+\sum_{k=1}^{n}|a_{k}(0)|^{2}|\sigma^{\prime}(w_{k}(0)x_{i}-b_{k}(0))|^{2}(x_{i}^{2}+1)\Big{)}$
$\leq m+\sum_{i=1}^{m}\sum_{k=1}^{n}\Big{(}2|w_{k}(0)|^{2}|x_{i}|^{2}+2|b_{k}(0)|^{2}+|a_{k}(0)|^{2}(|x_{i}|^{2}+1)\Big{)}$
$\leq m(1+2B_{x}^{2}\|\boldsymbol{w}(0)\|_{2}^{2}+2\|\boldsymbol{b}(0)\|_{2}^{2}+(B_{x}^{2}+1)n^{-1}\|\sqrt{n}\boldsymbol{a}(0)\|_{2}^{2}).$
Using a concentration inequality for sub-Gaussian random variables, we have $\|\boldsymbol{w}(0)\|_{2}$, $\|\boldsymbol{b}(0)\|_{2}$ and $\|\sqrt{n}\boldsymbol{a}(0)\|_{2}$ can be bounded by $C\sqrt{n}$ with some constant $C>0$. Therefore, by (38) we have $\|J_{\theta}\mathbf{f}(0)\|_{2}=O_{p}(\sqrt{n})$. Note a similar bound applies to $\|J_{\theta}f(x,\hat{\theta})\|_{2}$. Then for large enough $n$ we have
$\|H(0,x)\|_{2}\leq\frac{1}{n}\|J_{\theta}f(x,\hat{\theta})\|_{2}\cdot\frac{1}{\phi^{\prime\prime}(0)}\cdot\|J_{\theta}\mathbf{f}(0)\|_{2}=O_{p}(1).$
@@
$\leq\int_{0}^{t}\Big{|}\eta_{0}H(s,x)(\mathbf{f}(s)-\mathbf{y})-\eta_{0}\tilde{H}(s,x)(\tilde{\mathbf{f}}(s)-\mathbf{y})\Big{|}\mathrm{d}s$
$\leq\int_{0}^{t}\eta_{0}(\|H(0,x)\|_{2}+\|H(s,x)-H(0,x)\|_{2})\|\mathbf{r}(s)\|_{2}\mathrm{d}s$
$\quad+\int_{0}^{t}\eta_{0}\|H(s,x)-\tilde{H}(s,x)\|_{2}\|\tilde{\mathbf{f}}(s)-\mathbf{y}\|_{2}\mathrm{d}s$
@@ A.2 Implicit bias in parameter space
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\tilde{\boldsymbol{a}}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}[w_{k}(0)x_{i}-b_{k}(0)]_{+}=y_{i},\ \forall i\in[m].$ (39)
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)[wx-b]_{+}\ \mathrm{d}\mu_{n}\\ g(x,\alpha)=\int\alpha(w,b)[wx-b]_{+}\ \mathrm{d}\mu.\end{dcases}$ (40)
@@
Then given any $x\in\mathbb{R}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.
@@
[Proof of Proposition 16] Recall $\{w_{k}(0),b_{k}(0)\}_{k=1}^{n}$ denote the sampled initial input weights and biases, and $\mu_{n}$ denotes the associated empirical measure. We use the notation $\|\cdot\|_{n}$ for the $L_{2}(\mu_{n})$-seminorm in $\Theta_{\mathrm{ReLU}}$, i.e., for $\alpha\in\Theta_{\mathrm{ReLU}}$,
@@
Define $\sigma_{i}(w,b)=\sigma(wx_{i}-b)=[wx_{i}-b]_{+}\in\Theta_{\mathrm{ReLU}}$. Since $\bar{\boldsymbol{a}}$ is the solution to (39), we have that $\bar{\alpha}_{n}$ solves
@@
Note the objectives and constraints in problem (42) and problem (43) are essentially defined in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Specifically, for any $\alpha,\alpha^{\prime}\in\Theta_{\mathrm{ReLU}}$, if $\|\alpha-\alpha^{\prime}\|_{n}=0$, then $\alpha$ and $\alpha^{\prime}$ are equivalent with respect to feasibility and objective function value in both problems. Therefore, we shall address these two problems in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Note $L_{n}$ and $L_{n}^{\prime}$ are strictly convex in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Hence, $\bar{\alpha}_{n}$ and $\bar{\alpha}_{n}^{\prime}$ are unique in the $L_{2}(\mu_{n})$ sense.
@@
For a given data $x\in R$, notice that
$\|\sigma(wx-b)\|_{n}^{2}\leq\frac{1}{n}\sum_{k=1}^{n}|w_{k}(0)x-b_{k}(0)|^{2}\leq|x|+\frac{1}{n}\|\boldsymbol{b}(0)\|_{2}^{2}=O_{p}(1).$
@@
$\big{|}g_{n}(x,\bar{\alpha}_{n})-g_{n}(x,\bar{\alpha}_{n}^{\prime})\big{|}$ $=\big{|}\langle\bar{\alpha}_{n}(w,b)-\bar{\alpha}^{\prime}_{n}(w,b),\sigma(wx-b)\rangle_{n}\big{|}$ (47)
$\leq\|\bar{\alpha}_{n}-\bar{\alpha}^{\prime}_{n}\|_{n}\|\sigma(wx-b)\|_{n}$
@@
Given a fixed $x\in\mathbb{R}$, recall that
$\begin{dcases}g_{n}(x,\bar{\alpha})=\int\bar{\alpha}(w,b)[wx-b]_{+}\mathrm{d}\mu_{n};\\ g(x,\bar{\alpha})=\int\bar{\alpha}(w,b)[wx-b]_{+}\mathrm{d}\mu.\end{dcases}$
@@
$|g_{n}(x,\bar{\alpha})-g_{n}(x,\bar{\alpha}_{n}^{\prime})|$ $\leq\int\left(\sum_{i=1}^{m}|\lambda_{i}-\lambda_{i}^{(n)}|\sigma_{i}(w,b)\right)[wx-b]_{+}\mathrm{d}\mu_{n}$
$\leq\|\lambda-\lambda^{(n)}\|_{\infty}\int\left(\sum_{i=1}^{m}\sigma_{i}(w,b)\right)[wx-b]_{+}\mathrm{d}\mu_{n}.$
@@
Hence, with large enough $n$ and with high probability, $\int[wx-b]_{+}\sum_{j=1}^{m}\sigma_{j}(w,b)\ \mathrm{d}\mu_{n}$ can be bounded by a finite number independent of $n$. Therefore, by (50) we have
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)|wx-b|\ \mathrm{d}\mu_{n}\\ g(x,\alpha)=\int\alpha(w,b)|wx-b|\ \mathrm{d}\mu.\end{dcases}$ (52)
@@
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\tilde{\boldsymbol{a}}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}|w_{k}(0)x_{i}-b_{k}(0)|=y_{i},\ \forall i\in[m].$
@@
Then given any $x\in\mathbb{R}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.

### v2

@@ A.1 Linearization
In the following, we use the notation $\Delta=\sup_{t\geq 0}\|H(t)-\tilde{H}(t)\|_{2}$. Now we show $\boldsymbol{f}(t)$ remains close to $\tilde{\boldsymbol{f}}(t)$. Let $\boldsymbol{r}(t)=\boldsymbol{f}(t)-\tilde{\boldsymbol{f}}(t)$ and $C_{2}$ be the constant in Proposition 14. Consider the function $u(t)=\|\boldsymbol{r}(t)\|_{2}$ defined on $t\geq 0$. Clearly, $u(t)$ is continuous. Since $u(t)=0$ if and only if $\boldsymbol{r}(t)=0$, we have that $u(t)$ is differentiable whenever $u(t)\neq 0$. Note by Proposition 14, we have that, with high probability, $\lambda_{\min}(H(t))\geq C_{2}$ for all $t\geq 0$. Then when $u(t)\neq 0$, we have
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))^{2}$ $=\frac{\mathrm{d}}{\mathrm{dt}}\|e^{C_{2}\eta_{0}t}\boldsymbol{r}(t)\|_{2}^{2}$ (34)
$=2e^{C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}\boldsymbol{r}(t))$
$=2e^{2C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}\Big{(}C_{2}\eta_{0}\boldsymbol{r}(t)-\eta_{0}H(t)(\boldsymbol{f}(t)-\boldsymbol{y})+\eta_{0}\tilde{H}(t)(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})\Big{)}$
$=2\eta_{0}e^{2C_{2}\eta_{0}t}\Big{(}\boldsymbol{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})-\boldsymbol{r}(t)^{T}(H(t)-C_{2}I_{n})\boldsymbol{r}(t)\Big{)}$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}\boldsymbol{r}(t)^{T}(\tilde{H}(t)-H(t))(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y})$
$\leq 2\eta_{0}e^{2C_{2}\eta_{0}t}u(t)\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2}.$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}(e^{C_{2}\eta_{0}t}u(t))\leq\eta_{0}e^{C_{2}\eta_{0}t}\|\tilde{H}(t)-H(t)\|_{2}\|\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}\|_{2}.$
Now for any $t\geq 0$, consider $t^{\prime}=\inf\{s\in[0,t]\colon\boldsymbol{r}(s)\neq 0\}$. Then we have
$e^{C_{2}\eta_{0}t}u(t)\leq\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s.$
@@
$\|\boldsymbol{r}(t)\|_{2}$ $\leq e^{-C_{2}\eta_{0}t}\int_{t^{\prime}}^{t}\eta_{0}e^{C_{2}\eta_{0}s}\|\tilde{H}(s)-H(s)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s$ (35)
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}\Delta\int_{t^{\prime}}^{t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}\mathrm{d}s$
$\leq\eta_{0}e^{-C_{2}\eta_{0}t}\|\boldsymbol{f}(0)-\boldsymbol{y}\|_{2}t\Delta.$
Given $x\in\mathbb{R}^{d}$, by the chain rule we have:
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\theta(t))$ $=-\eta_{0}H(t,x)(\boldsymbol{f}(t)-\boldsymbol{y}),$ (36)
where $\text{where }H(t,x)$ $\triangleq\frac{1}{n}J_{\theta}f(x,\theta(t))(\nabla^{2}\Phi(\theta(t)))^{-1}J_{\theta}\boldsymbol{f}(t)^{T}$
@@
$\frac{\mathrm{d}}{\mathrm{dt}}f(x,\tilde{\theta}(t))$ $=-\eta_{0}\tilde{H}(t,x)(\tilde{\boldsymbol{f}}(t)-\boldsymbol{y}),$ (37)
where $\text{where }\tilde{H}(t,x)$ $\triangleq\frac{1}{n}J_{\boldsymbol{a}}f(x,\hat{\theta})(\nabla^{2}\Phi(\tilde{\boldsymbol{a}}(t)))^{-1}J_{\boldsymbol{a}}\boldsymbol{f}(0)^{T}.$
@@
$\|J_{\theta}\boldsymbol{f}(0)\|_{F}^{2}$ $=\sum_{i=1}^{m}\Big{(}1+\sum_{k=1}^{n}|\sigma(w_{k}(0)^{T}x_{i}-b_{k}(0))|^{2}$ (39)
$\quad+\sum_{k=1}^{n}|a_{k}(0)|^{2}|\sigma^{\prime}(w_{k}(0)^{T}x_{i}-b_{k}(0))|^{2}(\|x_{i}\|_{2}^{2}+1)\Big{)}$
$\leq m+\sum_{i=1}^{m}\sum_{k=1}^{n}\Big{(}2\|w_{k}(0)\|_{2}^{2}\|x_{i}\|_{2}^{2}+2|b_{k}(0)|^{2}+|a_{k}(0)|^{2}(\|x_{i}\|_{2}^{2}+1)\Big{)}$
$\leq m(1+2B_{x}^{2}+2\|\boldsymbol{b}(0)\|_{2}^{2}+(B_{x}^{2}+1)\|\boldsymbol{a}(0)\|_{2}^{2}).$
Using a concentration inequality for sub-Gaussian random variables, we have $\|\boldsymbol{b}(0)\|_{2}$ and $\|\sqrt{n}\boldsymbol{a}(0)\|_{2}$ can be bounded by $C\sqrt{n}$ with some constant $C>0$. Therefore, by (39) we have $\|J_{\theta}\boldsymbol{f}(0)\|_{2}=O_{p}(\sqrt{n})$. Note a similar bound applies to $\|J_{\theta}f(x,\hat{\theta})\|_{2}$. Then for large enough $n$ we have
$\|H(0,x)\|_{2}\leq\frac{1}{n}\|J_{\theta}f(x,\hat{\theta})\|_{2}\cdot\frac{1}{\phi^{\prime\prime}(0)}\cdot\|J_{\theta}\boldsymbol{f}(0)\|_{2}=O_{p}(1).$
@@
$\leq\int_{0}^{t}\Big{|}\eta_{0}H(s,x)(\boldsymbol{f}(s)-\boldsymbol{y})-\eta_{0}\tilde{H}(s,x)(\tilde{\boldsymbol{f}}(s)-\boldsymbol{y})\Big{|}\mathrm{d}s$
$\leq\int_{0}^{t}\eta_{0}(\|H(0,x)\|_{2}+\|H(s,x)-H(0,x)\|_{2})\|\boldsymbol{r}(s)\|_{2}\mathrm{d}s$
$\quad+\int_{0}^{t}\eta_{0}\|H(s,x)-\tilde{H}(s,x)\|_{2}\|\tilde{\boldsymbol{f}}(s)-\boldsymbol{y}\|_{2}\mathrm{d}s$
@@ A.2 Implicit bias in parameter space
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\tilde{\boldsymbol{a}}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}[w_{k}(0)^{T}x_{i}-b_{k}(0)]_{+}=y_{i},\ \forall i\in[m].$ (40)
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)[w^{T}x-b]_{+}\ \mathrm{d}\mu_{n}\\ g(x,\alpha)=\int\alpha(w,b)[w^{T}x-b]_{+}\ \mathrm{d}\mu.\end{dcases}$ (41)
@@
Then given any $x\in\mathbb{R}^{d}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.
@@
[Proof of Proposition 17] Recall $\{w_{k}(0),b_{k}(0)\}_{k=1}^{n}$ denote the sampled initial input weights and biases, and $\mu_{n}$ denotes the associated empirical measure. We use the notation $\|\cdot\|_{n}$ for the $L^{2}(\mu_{n})$-seminorm in $\Theta_{\mathrm{ReLU}}$, i.e., for $\alpha\in\Theta_{\mathrm{ReLU}}$,
@@
Define $\sigma_{i}(w,b)=[w^{T}x_{i}-b]_{+}\in\Theta_{\mathrm{ReLU}}$. Since $\bar{\boldsymbol{a}}$ is the solution to (40), we have that $\bar{\alpha}_{n}$ solves
@@
Note the objectives and constraints in problem (43) and problem (44) are essentially defined in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Specifically, for any $\alpha,\alpha^{\prime}\in\Theta_{\mathrm{ReLU}}$, if $\|\alpha-\alpha^{\prime}\|_{n}=0$, then $\alpha$ and $\alpha^{\prime}$ are equivalent with respect to feasibility and objective function value in both problems. Therefore, we shall address these two problems in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Note $L_{n}$ and $L_{n}^{\prime}$ are strictly convex in $\Theta_{\mathrm{ReLU}}/_{\|\cdot\|_{n}}$. Hence, $\bar{\alpha}_{n}$ and $\bar{\alpha}_{n}^{\prime}$ are unique in the $L^{2}(\mu_{n})$ sense.
@@
For a given data $x\in\mathbb{R}^{d}$, notice that
$\|\sigma(w^{T}x-b)\|_{n}^{2}\leq\frac{1}{n}\sum_{k=1}^{n}|w_{k}(0)^{T}x-b_{k}(0)|^{2}\leq\|x\|+\frac{1}{n}\|\boldsymbol{b}(0)\|_{2}^{2}=O_{p}(1).$
@@
$\big{|}g_{n}(x,\bar{\alpha}_{n})-g_{n}(x,\bar{\alpha}_{n}^{\prime})\big{|}$ $=\big{|}\langle\bar{\alpha}_{n}(w,b)-\bar{\alpha}^{\prime}_{n}(w,b),\sigma(w^{T}x-b)\rangle_{n}\big{|}$ (48)
$\leq\|\bar{\alpha}_{n}-\bar{\alpha}^{\prime}_{n}\|_{n}\|\sigma(w^{T}x-b)\|_{n}$
@@
Given a fixed $x\in\mathbb{R}^{d}$, recall that
$\begin{dcases}g_{n}(x,\bar{\alpha})=\int\bar{\alpha}(w,b)[w^{T}x-b]_{+}\mathrm{d}\mu_{n};\\ g(x,\bar{\alpha})=\int\bar{\alpha}(w,b)[w^{T}x-b]_{+}\mathrm{d}\mu.\end{dcases}$
@@
$|g_{n}(x,\bar{\alpha})-g_{n}(x,\bar{\alpha}_{n}^{\prime})|$ $\leq\int\left(\sum_{i=1}^{m}|\lambda_{i}-\lambda_{i}^{(n)}|\sigma_{i}(w,b)\right)[w^{T}x-b]_{+}\mathrm{d}\mu_{n}$
$\leq\|\lambda-\lambda^{(n)}\|_{\infty}\int\left(\sum_{i=1}^{m}\sigma_{i}(w,b)\right)[w^{T}x-b]_{+}\mathrm{d}\mu_{n}.$
@@
Hence, with large enough $n$ and with high probability, $\int[w^{T}x-b]_{+}\sum_{j=1}^{m}\sigma_{j}(w,b)\ \mathrm{d}\mu_{n}$ can be bounded by a finite number independent of $n$. Therefore, by (51) we have
@@
$\begin{dcases}g_{n}(x,\alpha)=\int\alpha(w,b)|w^{T}x-b|\ \mathrm{d}\mu_{n}\\ g(x,\alpha)=\int\alpha(w,b)|w^{T}x-b|\ \mathrm{d}\mu.\end{dcases}$ (53)
@@
$\min_{\tilde{\boldsymbol{a}}\in\mathbb{R}^{n}}\ D_{\Phi}(\tilde{\boldsymbol{a}},\tilde{\boldsymbol{a}}(0))\quad\mathrm{s.t.}\ \sum_{k\in[n]}\tilde{a}_{k}|w_{k}(0)^{T}x_{i}-b_{k}(0)|=y_{i},\ \forall i\in[m].$
@@
Then given any $x\in\mathbb{R}^{d}$, with high probability over initialization, $\lim_{n\rightarrow\infty}|g_{n}(x,\bar{\alpha}_{n})-g(x,\bar{\alpha})|=0$.